The volume of this cone is 471 ft^3 and the height is 18 ft. I need to find the radius.
Answer by rfer(16322)

V=Pir^2h/3
471=(3.14*r^2*18)/3
471=56.52r^2/3
471*3=56.52r^2
1413=56.52r^2
1413/56.52=r^2
25=r^2
5=r